OpenStudy (anonymous):

A non-uniform lever of mass 14.7 kg and of length 3 m is pivoted at its mid-point. The lever will be in equilibrium when an object of mass 6.7 kg hangs from its right end. Find the location of its center of gravity?

OpenStudy (anonymous):

if it gets to equilibrium when second mass hangs from right end, it means center of mas needs to be on the left. \[14.7*x=6.7*(3/2)\] x=0.68 to left from center